OpenStudy (amilapsn):

Show that r! | n(n+1)...(n+r-1) Here n,r are positive integers.....

OpenStudy (amilapsn):

@Marki

OpenStudy (anonymous):

lol no its ok i miss read

OpenStudy (amilapsn):

ok

OpenStudy (anonymous):

familiar with residues\mod ?

OpenStudy (amilapsn):

no i havent heard about them...

OpenStudy (anonymous):

ok remainder ?

OpenStudy (amilapsn):

yes

OpenStudy (amilapsn):

whaen we divide 12 by 5 remainder is 2 ok?

OpenStudy (anonymous):

lets assume n leaves 1 as remainder when divided by r (without lose of generality )

OpenStudy (amilapsn):

ok

OpenStudy (anonymous):

so n leaves 1 as remainder n+1 leaves 2 as remainder n+3 leaves 3 as remainder . . . n+r-1 leaves r-1 as remainder agree ?

OpenStudy (amilapsn):

yes...

OpenStudy (anonymous):

wait sorry typo at last step so n leaves 1 as remainder n+1 leaves 2 as remainder n+3 leaves 3 as remainder . . . n+r-1 leaves r as remainder

OpenStudy (amilapsn):

r as remainder means no remainder.... right?

OpenStudy (anonymous):

now u can assume n have r remainder or anything u want ... this is what it means W.L.O.G

OpenStudy (anonymous):

yes but also means it divides r , so we can write it as r instead of 0

OpenStudy (amilapsn):

ok

OpenStudy (amilapsn):

hey let me prove...!

OpenStudy (anonymous):

now n(n+1)...(n+r-1) is equivalent to 1.2.3.4......r (multiple of remainders , right )

OpenStudy (amilapsn):

Thank you so much...

OpenStudy (anonymous):

ok go

OpenStudy (amilapsn):

n mod r=r n+1 mod r=r+1 ....... (n+r-1) mod r=2r-1

OpenStudy (amilapsn):

oh no i think i messed up... @Marki

OpenStudy (anonymous):

hmm n+r-1=r-1 mod r

OpenStudy (anonymous):

so it would be equivalent to 1.2.3....(r-1) r mod r , (r)! which divide r!

OpenStudy (amilapsn):

how can we say that n(n+1)...(n+r-1) is equivalent to 1.2.3...r?

OpenStudy (anonymous):

like this \(n(n+1)...(n+r-1) \equiv 1.2.3.....r \mod r\)

OpenStudy (amilapsn):

sorry :-( i didnt get that....

OpenStudy (thomas5267):

\(n<r\)? Clearly the thing won't work if n=10 and r=2...

OpenStudy (anonymous):

no it works :)

OpenStudy (anonymous):

n=10 , n+1=11,n+2=12|2!

OpenStudy (thomas5267):

Did I read it backwards? Are you looking for n(n+1)(n+2)...(n+r-1) divides r! or r! divides n(n+1)(n+2)...(n+r-1)?

OpenStudy (anonymous):

haha my bad xD

OpenStudy (thomas5267):

Something is wrong here. Let n=10, r=2 10*11=110 2!=2 Clearly 110 does not divide 2.

OpenStudy (amilapsn):

Is my question wrong?

ganeshie8 (ganeshie8):

i love wilson theorem xD i also like below combinatoric proof as it looks short and neat\[\binom{n}{r} = \dfrac{n!}{(n-r)!r!} \in \mathbb{Z}\]

OpenStudy (anonymous):

i was proving r!| n(n+1)...(n+r-1)

ganeshie8 (ganeshie8):

yes ^

ganeshie8 (ganeshie8):

i have edited the question, see if it looks good now :)

OpenStudy (thomas5267):

One of my textbook wrote a divides b as b|a so I am not surprised that there are some notation confusion.

ganeshie8 (ganeshie8):

\(a ~|~ b\) is read as \(a\) divides \(b\)

ganeshie8 (ganeshie8):

examples : 2 | 6 5 | 25 7 | 28 @amilapsn

OpenStudy (thomas5267):

Yeah my textbook is full of errors.

OpenStudy (amilapsn):

ok my bad.... ><

ganeshie8 (ganeshie8):

No, it was a good catch @thomas5267 :) i overlooked that mistake in the question earlier haha

OpenStudy (thomas5267):

Is it a proof by induction?

ganeshie8 (ganeshie8):

refering to combinotoric proof, @thomas5267 ?

OpenStudy (thomas5267):

I don't understand your proof to be honest. Could you elaborate?

ganeshie8 (ganeshie8):

sure, do we agree that the binomial coefficient is an integer because it simply represents the number of ways of choosing "r" different things from a set of "n" things ?

ganeshie8 (ganeshie8):

\[\binom{n}{r} = \dfrac{n!}{(n-r)!r!}\] is an integer because this represents the "count" of number of ways of choosing "r" things from "n" things

OpenStudy (thomas5267):

But OP is asking for r!|n(n+1)(n+2)...(n+r-1) not r!|n(n-1)(n-2)...(n-r+1)...

OpenStudy (anonymous):

wait does that mean my solution was right ?

ganeshie8 (ganeshie8):

Right, see below : \[n(n+1)(n+2)\ldots (n-r+1) = \dfrac{n(n+1)(n+2)\ldots (n-r+1)\color{red}{(n-r)(n-r-1)\ldots 3 .2. 1}}{\color{red}{(n-r)(n-r-1)\ldots 3 .2. 1}} \]

ganeshie8 (ganeshie8):

\[ = \dfrac{n!}{(n-r)!}\]

ganeshie8 (ganeshie8):

from the binomial coefficient you know that \(\dfrac{n!}{(n-r)!r!}\) is an integer that means \(r! ~\Bigg| ~\dfrac{n!}{(n-r)!}\)

OpenStudy (anonymous):

its nice one

ganeshie8 (ganeshie8):

yes marki, your proof looks good to me !

OpenStudy (anonymous):

haha no im saying its nice one about urs >.<

OpenStudy (thomas5267):

I am totally confused lol.

ganeshie8 (ganeshie8):

let me break this for you into 2-3 steps :)

ganeshie8 (ganeshie8):

do you agree that `n(n+1)...(n+r-1)` can be written as `n!/(n-r)!` ?

OpenStudy (thomas5267):

I don't get this part.

ganeshie8 (ganeshie8):

I see.. remember the permutation formula ? nPr ?

ganeshie8 (ganeshie8):

lets prove this ` n(n+1)...(n+r-1) = n!/(n-r)!`a

ganeshie8 (ganeshie8):

lets start with the right hand side : \[\dfrac{n!}{(n-r)!}\]

ganeshie8 (ganeshie8):

expand the factorial down to n-r+1 term : \[\dfrac{n!}{(n-r)!} = \dfrac{n(n-1)(n-2)\cdots (n-r+1)(n-r)!}{(n-r)!}\]

ganeshie8 (ganeshie8):

we're done! just cancel the (n-r)! top and bottom

ganeshie8 (ganeshie8):

for example : \[\dfrac{10!}{(10-3)!} = \dfrac{10(10-1)(10-2)(10-3)!}{(10-3)!} = 10(10-1)(10-2)\]

OpenStudy (anonymous):

(n+r-1)!=1.2.3.....(n-1)n(n+1)....(n+r-1) n(n+1).....(n+r-1)=(n+r-1)!/(n-1)! hehehe

ganeshie8 (ganeshie8):

another example : \[\dfrac{14!}{12!} = \dfrac{14\cdot 13\cdot 12!}{12!} = 14\cdot 13 \]

OpenStudy (thomas5267):

Yes but how does n(n+1)(n+2)...(n+r-1)=n(n-1)(n-2)(n-3)...(n-r+1)?

ganeshie8 (ganeshie8):

Oops! i see what you're asking now :|

OpenStudy (anonymous):

n!/(n-r)!=1.2.3..(n-r)(n-r+1)...n/1.2.3....(n-r) =(n-r+1)(n-r+2)....n hmm

OpenStudy (thomas5267):

I am too tired to continue. I have to go eat my dinner. Cya.

OpenStudy (anonymous):

n!/(n-r)!=n(n-1)(n-2)....(n-r+2)(n-r+1)

OpenStudy (anonymous):

i agree

ganeshie8 (ganeshie8):

it doesnt matter however, just consider (n+r-1)!/(n-1)! and argue the same